OpenStudy (anonymous):

if p is a polynomial how do i show that the limit of p(x) as x approaches a = p(a)

OpenStudy (anonymous):

\[\lim_{x \rightarrow a}p(x)=p(a)\]

OpenStudy (fibonaccichick666):

well... i need some more info first

OpenStudy (fibonaccichick666):

what type of class is this for, analysis or calc?

OpenStudy (anonymous):

calc

OpenStudy (fibonaccichick666):

oook, uhm well, crap... hmm what arsenal do you have so far? Have you done the addition of limits?

OpenStudy (anonymous):

yeah

OpenStudy (fibonaccichick666):

ok so can you tell me what every single polynomial looks like?

OpenStudy (fibonaccichick666):

the general form of a polynomial

OpenStudy (anonymous):

x^2+x+c

OpenStudy (fibonaccichick666):

well, almost what about a 5th deg poly? or a 10th degree?

OpenStudy (anonymous):

x^n+x^n-1+...+c

OpenStudy (fibonaccichick666):

how about coefficients too?

OpenStudy (anonymous):

anx^n+an-1x^(n-1)+an-2x^(n-2)+...a0

OpenStudy (fibonaccichick666):

good, ok, so now, what if you took the limit of that

OpenStudy (fibonaccichick666):

could you split it up?

OpenStudy (anonymous):

wouldnt you just get a's instead of x's

OpenStudy (fibonaccichick666):

no, not quite, what happens is you change it into the sum of limits, which hopefully you have said \(lim_{x\rightarrow c} x^2=c^2\)

OpenStudy (fibonaccichick666):

and similar, so you can just apply that then undo in the end

OpenStudy (fibonaccichick666):

So we want \[\lim_{x\rightarrow c} a_nx^n+a_{n-1}x^{(n-1)}+a_{{n-2}}x^{(n-2)}+...+a_0=?\]

OpenStudy (fibonaccichick666):

We can use addition there right to split up the limit?

OpenStudy (anonymous):

yeah \[\lim_{x \rightarrow c}a _{n}x ^{n}+\lim_{x \rightarrow c} a _{n-1}x ^{(n-1)}+....\] like this right?

OpenStudy (fibonaccichick666):

good, now, have you shown in class what \[lim_{x\rightarrow c} x^n=?\]

OpenStudy (anonymous):

isn't it just \[c^n\]

OpenStudy (fibonaccichick666):

it is, but if you haven't shown it, we can't use it :/

OpenStudy (anonymous):

oh how do you show it

OpenStudy (fibonaccichick666):

well, have you done product of limits? or have you learned the epsilon delta definition of a limit?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

you haven't done product of a limit is the limit of the product?

OpenStudy (anonymous):

i've done the product of limits but not epsilon delta proof

OpenStudy (fibonaccichick666):

ok that's fine then, so now, how else can you write x^n?

OpenStudy (anonymous):

x multiplied by n number of x's

OpenStudy (fibonaccichick666):

good good, ok so limit of x?

OpenStudy (fibonaccichick666):

as x goes to c, sorry forgot to mention

OpenStudy (anonymous):

c?

OpenStudy (fibonaccichick666):

yes, and I'm really hoping you've covered that or else based off of info given, you cannot do this problem

OpenStudy (fibonaccichick666):

i guess you could just say it, it is a definition after all

OpenStudy (anonymous):

oh ok, we weren't really shown in depth on the properties of limits and stuff, so i'm a bit lost

OpenStudy (fibonaccichick666):

well, using \(\epsilon-\delta\) you can prove it but you just have to accept that

OpenStudy (fibonaccichick666):

oh ok, so. Let's say the only things you know are that lim x=c and lim k=k if k is a constant

OpenStudy (fibonaccichick666):

So, a limit can be distributed and factored but you must remember it is a function so it cannot be divided away

OpenStudy (fibonaccichick666):

so you have to use a 2 step proof here. First, you need to show that lim ax^n=ac^n, then use that to show your big lim p(x)=p(c)

OpenStudy (fibonaccichick666):

if you would like a better explanation of the limit properties you can look on khan academy and purplemath

OpenStudy (anonymous):

but i can't show that lim ax^n=ac^n yet so i just leave it at that?

OpenStudy (fibonaccichick666):

yea, you can... hmm. give me a min, read this for me http://tutorial.math.lamar.edu/Classes/CalcI/LimitsProperties.aspx

OpenStudy (fibonaccichick666):

our goal will be to ONLY use those properties in order to manipulate the polynomial into something we can do

OpenStudy (fibonaccichick666):

And we will be able to assume lim x=c and lim k=k

OpenStudy (fibonaccichick666):

but that is it

OpenStudy (anonymous):

i see ty for the help

OpenStudy (fibonaccichick666):

np, if you want to try writing that up, I can check it, but this is something where you may not have enough info in class to answer

OpenStudy (fibonaccichick666):

It all depends on those properties I linked to